OpenStudy (lgbasallote):

the gravitational force exerted on a baseball is 2.21 N down, a pitcher throws the ball horizontally with velocity 18.0 m/s by uniformly accelerating it along a straight horizontal line for a time interval of 170 ms. the ball starts from rest. Through what distance does it move before its release? what are the magnitude and direction of the force of the pitcher exerts on the ball?

OpenStudy (anonymous):

what's the unit for that time interval?

OpenStudy (lgbasallote):

milisecond i suppose

OpenStudy (anonymous):

i got the magnitude of F = 23.88N

OpenStudy (lgbasallote):

explain?

OpenStudy (anonymous):

gravitational force acting on the ball is F = mg = 2.21 N........from this we can get m. then we've been given, v = 18 m/s, u = 0, t = 170 * 10^(-3) sec. sub in v = U + at...........this way we can get a. and F = ma.

OpenStudy (lgbasallote):

hmm

OpenStudy (lgbasallote):

but it's looking for distance right?

OpenStudy (anonymous):

i know, i started with the second part pf the question. it seemed easy!

OpenStudy (lgbasallote):

hmm okay..

OpenStudy (lgbasallote):

i got 106 m/s^2

OpenStudy (anonymous):

yea, me too.

OpenStudy (lgbasallote):

i got F = 23.88 N

OpenStudy (lgbasallote):

so that's the magnitude?

OpenStudy (anonymous):

i think so. and its direction should be horizontal because that's the direction of acceleration.

OpenStudy (lgbasallote):

but positive or negative?

OpenStudy (anonymous):

positive.

OpenStudy (lgbasallote):

hmmm so what's the first question

OpenStudy (anonymous):

still thinking about that...

OpenStudy (anonymous):

do u have any way of confirming the answers?

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

I'm not too sure abt this: should the distance traveled before its release be 0? we know W = F.s also W = change in KE = 1/2 m (v^2 - u^2) but the question says "the ball starts from rest". so the change in velocity = 0 thus W = 0. hence, s = 0. what do u think?

OpenStudy (anonymous):

is the motion look likes this?